The sum of the digits of a certain two-digit number is 7. Reversing its digits increases the number by 9.
What is the number?

Answer by edjones(8007)

Let u=the units digit and t=the tens digit. The digit = 10t+u
t+u=7
.
10t+u=10u+t-9
9t-9u=-9
9(t-u)=-9
t-u=-1
.
t+u=7
t-u=-1
-------Add
2t=6
.
t=3
u=4
The number is 34.
.
Ed